### v1

@@ Abstract
The advent of vision-language models fosters the interactive conversations between AI-enabled models and humans. Yet applying these models into clinics must deal with daunting challenges around large-scale training data, financial, and computational resources. Here we propose a cost-effective instruction learning framework for conversational pathology named as CLOVER. CLOVER only trains a lightweight module and uses instruction tuning while freezing the parameters of the large language model. Instead of using costly GPT-4, we propose well-designed prompts on GPT-3.5 for building generation-based instructions, emphasizing the utility of pathological knowledge derived from the Internet source. To augment the use of instructions, we construct a high-quality set of template-based instructions in the context of digital pathology. From two benchmark datasets, our findings reveal the strength of hybrid-form instructions in the visual question-answer in pathology. Extensive results show the cost-effectiveness of CLOVER in answering both open-ended and closed-ended questions, where CLOVER outperforms strong baselines that possess 37 times more training parameters and use instruction data generated from GPT-4. Through the instruction tuning, CLOVER exhibits robustness of few-shot learning in the external clinical dataset. These findings demonstrate that cost-effective modeling of CLOVER could accelerate the adoption of rapid conversational applications in the landscape of digital pathology.
keywords:
Vision-language model, Large language model (LLM), Generative AI, Digital pathology, Prompt engineering
1 Main
@@
Emerging pathological vision-language models (PVLMs) (e.g., LLaVA-Med [18] and Quilt-LLaVA [19]) have demonstrated their utility in analyzing medical imaging. However, it is widely known that building a capable PVLM demands an excessive training data, human labour, financial, and computational resources [20, 21]. Extending general-purpose models into pathology-oriented model starts with using pathological vision-language datasets. These datasets consist of (i) image-text pairs (image-caption) dataset and (ii) instruction (image-question-answering) dataset [22, 23, 24, 25]. The image-text dataset aligns visual and language features [26], providing rich semantic information for pathological image contents [27, 28, 29]. Meanwhile, instruction dataset is crucial for activating LLMs to complete the visual-language question answering. Yet the process of instruction generation incurs a considerable financial cost by using GPT-4 (nearly \$9,000) [19]. In addition to the instruction data, model optimization also requires a substantial compute support. Directly tuning a billion-parameter LLM demands high computational resources and is impractical to achieve on consumer-grade GPUs. To overcome these bottlenecks, fundamental questions are urged to be addressed towards building an effective but low-cost PVLM: (i) How can we come up with a lightweight tuning approach to complement the LLaVA-like tuning on large training parameters? (ii) How can we generate a useful instruction dataset in a cost-effective way? (iii) How can we achieve few-shot generalized learning ability of PVLM for clinical applications without using a large-scale instruction dataset?
2 Building Efficient PVLM via CLOVER
2.1 Vision-language Foundation Model
2.2 Medical Vision-language Foundation Model
Building medical vision-language foundation models requires systematic efforts in both data curation and model inference in the healthcare system [1]. The large-scale resources from the literature (PMC-15M [22]), social media (PLIP [24] and Quilt-1M [23]), accessible textbooks [21] and private hospital data [25, 34] provide critical data support for domain-specific multi-modal models. Examples include PMC-LLaMA [20], LLaVA-Med [18], and Med-Flamingo [21]. LLaVA-Med is a representative work that aligns vision and language using 600K image-text pairs on the basis of LLaVA and fine-tunes the model using 60K dialogue-based instructions. Yet these instruction sets are generated by GPT-4 at a high cost. Similarly, in the pathological field, Quilt-LLaVA [19] uses Quilt-1M and GPT-4 to generate instruction data at a cost of up to \$8,858. Towards a better utility of foundation models, instruction preparation and generation in the various forms of structure and scale require substantial research efforts. Meanwhile, during the fine-tuning stage, LLaVA-like methods [9] update the parameters of the LLM, resulting in a high training cost that can not be accommodated on consumer-grade GPUs. This pressing demand for high-performance training equipment becomes a daunting hurdle for researchers to extend general-purpose models into domain-specific vision-language models.
2.3 CLOVER
Figure 1: The workflow of CLOVER. CLOVER employs the training framework of BLIP-2 to achieve a fast domain tuning with lightweight parameters. The entire training process of CLOVER includes two major stages: (i) alignment of vision and language and (ii) supervised fine-tuning with instructions. The alignment compels the model to acquire valuable representations between vision and language. Instruction fine-tuning is vital here for activating LLMs to excel in visual language question answering. Stage 1 requires inputs of image-text pairs, where we use the large-scale Quilt-1M [23] dataset. Stage 2 demands domain-specific instruction data. As we have seen a significant lack of the required instruction data in the literature, we propose a low-cost solution of instruction data generation carefully designed for analyzing pathological data.
In this study, we propose a cost-effective learning framework for accurate pathology vision and language inference named as CLOVER (Fig. 1). To achieve fast domain tuning with low training resources, we adopt the BLIP-2 architecture [33] as a visual language pre-training using a lightweight trainable query transformer (Q-former), a frozen visual encoder, and a frozen LLM. We leverage the paired pathological image and text captures from the Quilt-1M [23] dataset to align the vision and language. For the instruction fine-tuning, we meticulously design prompts tailored for pathological question answering and use GPT-3.5 [7] for generating effective instruction dataset at low cost, referred as the “generation-based instructions”. The instructions generated by our PVLM-orient prompt show strong capabilities in the model tuning. Additionally, we construct instructions by matching template questions with the original text captions to enhance the model’s vision understanding ability, referred as the “template-based instructions”. Integrating these two types of instructions gives rise to the hybrid-form instructions, which remarkably enhance CLOVER’s conversational abilities in pathology. Using two benchmark datasets and one independent clinical dataset, we comprehensively validate the effectiveness of CLOVER and demonstrate its potential for assisting clinical pathological tasks by “talking to your pathology data” in resource-constrained settings.
@@ 2.3 CLOVER
Our study has made multifaceted contributions: First, we propose an effective PVLM training framework particularly at low computational resources and financial spending. Different from the LLaVA-like tuning (e.g. LLaVA-Med), we emphasize the use of BLIP-2 [33] to serve as an alternative choice for the cost-effective lightweight inference. We find that tuning the wide spectrum of LLM’s parameters is unnecessary for building a sufficiently usable PVLM. Second, the value of our designed PVLM-oriented prompt is pronounced on GPT-3.5 instead of the advanced GPT-4. Low-cost CLOVER model outperforms those models trained with LLM tuning on the instruction data generated by GPT-4 in multiple settings. We also recognize template-based instructions without relying on GPT can significantly supplement generation-based instructions. A combined use of these two type instructions can further enhance CLOVER’s understanding capabilities. Third, we confirm that CLOVER can be effectively developed with a limited scale of instructions towards a frugal and real-world application. This is evidenced by the appealing performance in fine-tuning and zero-shot experiments on two visual question answering datasets, as well as in the few-shot experiment on a real-world clinical dataset.
@@ 3.1 Quantitative Comparison
We systematically measure CLOVER’s visual question answering (VQA) performance using two public benchmark datasets, namely PathVQA [35] and QUILT-VQA [19], and one independent clinical data set (see Sec. Datasets). We compare CLOVER’s performance with state-of-the-art (SOTA) PVLMs trained with notable costs to observe the cost-effectiveness of CLOVER.
Comparison with SOTA PVLMs on PathVQA. From Table 1, CLOVER outperforms major competing methods, showing a remarkable improvement of 26.13% in accuracy at maximum for closed-ended questions. Our model performance is even approaching to LLaVA-Med (37 times more model parameters) with a minor accuracy difference of 1.83%. Note that LLaVA and LLaVA-Med achieve their results through extensive parameter tuning of LLM and training on a vast instruction dataset generated by GPT-4. To illustrate, our method involves training with a frozen LLM, adjusting only a small portion of parameters (about 1/37 of LLaVA-Med). Meanwhile, CLOVER only involves the use an entry-level GPT-3.5 with a smaller scale of instruction dataset (2/3 of LLaVA-Med). As seen in Table 1, our model also shows superior performance improvement in open-ended question scenarios. CLOVER’s performance is twice that of BLIP-2 and approaches five times that of LLaVA in open-ended question scenarios. Compared with BLIP-2, our model improves in both closed-ended and open-ended question-answering. This key finding suggests that vision-language models could evidently benefit from the high-quality pathology-sensitive instruction data towards a low-cost model development.
@@
Table 1: Comparison with SOTA methods on PathVQA dataset.
Comparison with SOTA PVLMs on QUILT-VQA. We validate CLOVER’s zero-shot generalization capability on the QUILT-VQA dataset. From Table 2, CLOVER achieves the leading performance in precision and F1-score. Our results are even close to LLaVA and LLaVA-Med in terms of recall. Due to the longer answer length of LLaVA and LLaVA-Med, these models are advantageous in the recall evaluation metric. The average length (word counts) of true answers is 18 and that number for LLaVA and LLaVA-Med are 38 and 36 respectively, both of which are higher than BLIP-2 and our model (4 and 22 respectively). In addition, our precision exceeds LLaVA’s by 9.77% and LLaVA-Med’s by 10.25%, while our F1 score also surpasses LLaVA’s by 6.73% and LLaVA-Med’s by 6.53%. In terms of a performance-cost ratio, CLOVER achieves the leading values for all metrics. Overall, these findings indicate the strength of CLOVER to retain the high-level performance while minimizing costs.

### v2

@@ Abstract
The advent of vision-language models fosters the interactive conversations between AI-enabled models and humans. Yet applying these models into clinics must deal with challenges around large-scale training data, financial, and computational resources. Here we propose CLOVER, a cost-effective instruction learning framework for conversational pathology. CLOVER trains a lightweight module and uses instruction tuning while freezing the parameters of the large language model. Instead of using costly GPT-4, we propose well-designed prompts on GPT-3.5 for building generation-based instructions, emphasizing the utility of pathological knowledge derived from the Internet source. We construct a high-quality set of template-based instructions in the context of digital pathology. From two benchmark datasets, our findings reveal the strength of hybrid-form instructions in pathological visual question-answer. CLOVER outperforms baselines that possess 37 times more training parameters and exhibits few-shot capacity in the external clinical dataset. CLOVER could thus accelerate the adoption of rapid conversational applications in digital pathology.
1 Introduction
@@
Emerging pathological vision-language models (PVLMs) (for instance, LLaVA-Med [18] and Quilt-LLaVA [19]) have demonstrated their utility in analyzing medical imaging. However, it is widely known that building a capable PVLM demands an excessive training data, human labour, financial, and computational resources [20, 21, 22, 23, 24, 25]. Extending general-purpose models into pathology-oriented model starts with using pathological vision-language datasets. These datasets consist of (i) image-text pairs (image-caption) dataset and (ii) instruction (image-question-answering) dataset [26, 27, 28, 23]. The image-text dataset aligns visual and language features [29], providing rich semantic information for pathological image contents [30, 31, 32]. Meanwhile, instruction dataset is crucial for activating LLMs to complete the visual-language question answering. Yet the process of instruction generation incurs a considerable financial cost by using GPT-4 (nearly \$9,000) [19]. In addition to the instruction data, model optimization also requires a substantial compute support. Directly tuning a billion-parameter LLM demands high computational resources and is impractical to achieve on consumer-grade GPUs. To overcome these bottlenecks, fundamental questions are urged to be addressed towards building an effective but low-cost PVLM: (i) How can we come up with a lightweight tuning approach to complement the LLaVA-like tuning on large training parameters? (ii) How can we generate a useful instruction dataset in a cost-effective way? (iii) How can we achieve few-shot generalized learning ability of PVLM for clinical applications without using a large-scale instruction dataset?
@@
In this study, to address these questions, we propose a cost-effective learning framework for accurate pathology vision and language inference named as CLOVER. Our study has made multifaceted contributions: First, we propose an effective PVLM training framework particularly at low computational resources and financial spending. Different from the LLaVA-like tuning (for instance, LLaVA-Med), we emphasize the use of BLIP-2 [33] to serve as an alternative choice for the cost-effective lightweight inference. We find that tuning the wide spectrum of LLM’s parameters is unnecessary for building a sufficiently usable PVLM. Second, the value of our designed PVLM-oriented prompt is pronounced on GPT-3.5 instead of the advanced GPT-4. Low-cost CLOVER model outperforms those models trained with LLM tuning on the instruction data generated by GPT-4 in multiple settings. We also recognize template-based instructions without relying on GPT can supplement generation-based instructions. A combined use of these two type instructions can further enhance CLOVER’s understanding capabilities. Third, we confirm that CLOVER can be effectively developed with a limited scale of instructions towards a frugal and real-world application. This is evidenced by the appealing performance in fine-tuning and zero-shot experiments on two visual question answering datasets, as well as in the few-shot experiment on a real-world clinical dataset.
Figure 1: (a) The workflow of instruction generation. We propose a low-cost solution of instruction data generation carefully designed for analyzing pathological data. Generation-based instructions are created by rewriting image captions into a question-and-answer format using low-cost GPT-3.5. In template-based instructions, the question is a descriptive prompt, while the answer is directly the image caption (more details in Method section 4.1). (b) The distribution of covered body parts or cancer types in our constructed instruction data. (c) The distribution of question-and-answer sentence lengths in our instruction data. (d) The workflow of CLOVER. CLOVER employs the training framework of BLIP-2 to achieve a fast domain tuning with lightweight parameters. The entire training process of CLOVER includes two major stages: (i) alignment of vision and language and (ii) supervised fine-tuning with instructions. The alignment compels the model to acquire valuable representations between vision and language. Instruction fine-tuning is vital here for activating LLMs to excel in visual language question answering. Stage 1 requires inputs of image-text pairs, where we use the large-scale Quilt-1M dataset. Stage 2 demands our self-constructed domain-specific instruction data.
@@ 2 Results
2.1 Model Overview
A schematic illustration of the CLOVER is offered in Fig. 1. To achieve fast domain tuning with low training resources, we adopt the BLIP-2 architecture [33] as a visual language pre-training using a lightweight trainable query transformer (Q-former), a frozen visual encoder, and a frozen LLM. We leverage the paired pathological image and text captures from the Quilt-1M [27] dataset to align the vision and language. For the instruction fine-tuning, we propose two approaches to generation instruction data (Fig. 1(a)). We meticulously design prompts tailored for pathological question answering and use GPT-3.5 [7] for generating effective instruction dataset at a low cost, referred as the “generation-based instructions”. The instructions generated by our PVLM-orient prompt show strong capabilities in the model tuning. Additionally, we construct instructions by matching template questions with the original text captions to enhance the model’s vision understanding ability, referred as the “template-based instructions”. See more details in Methods section. Integrating these two types of instructions gives rise to the hybrid-form instructions, which remarkably enhance CLOVER’s conversational abilities in pathology. Using two benchmark datasets and one independent clinical dataset, we comprehensively validate the effectiveness of CLOVER and demonstrate its potential for assisting clinical pathological tasks by “talking to your pathology data” in resource-constrained settings. We discuss the related literature to CLOVER in Supplementary Related Works section ).
@@ 2.2 Quantitative Comparison
We systematically measure CLOVER’s VQA performance using two public benchmark datasets, namely PathVQA [34] and QUILT-VQA [19], and one independent clinical data set (see Sec. Datasets). We compare CLOVER’s performance with state-of-the-art (SOTA) PVLMs trained with notable costs to observe the cost-effectiveness of CLOVER (see Sec. Compared Methods and Instructions for more details about other PVLMs).
From Table 1, CLOVER outperforms major competing methods on PathVQA dataset, showing a remarkable improvement of 26.13% in accuracy at maximum for closed-ended questions. Our model performance is even approaching to LLaVA-Med (37 times more model parameters) with a minor accuracy difference of 1.83%. Note that LLaVA and LLaVA-Med achieve their results through extensive parameter tuning of LLM and training on a vast instruction dataset generated by GPT-4. To illustrate, our method involves training with a frozen LLM, adjusting only a small portion of parameters (about 1/37 of LLaVA-Med). Meanwhile, CLOVER only involves the use an entry-level GPT-3.5 with a smaller scale of instruction dataset (2/3 of LLaVA-Med). As seen in Table 1, our model also shows superior performance improvement in open-ended question scenarios. CLOVER’s performance is twice that of BLIP-2 and approaches five times that of LLaVA in open-ended question scenarios. Compared with BLIP-2, our model improves in both closed-ended and open-ended question-answering. This key finding suggests that vision-language models could evidently benefit from the high-quality pathology-sensitive instruction data towards a low-cost model development.
We validate CLOVER’s zero-shot generalization capability on the QUILT-VQA dataset. From Table 2, CLOVER based on GPT-3.5 (default setting) achieves the leading performance in precision and F1-score. Our results are even close to LLaVA-Med and Quilt-LLaVA in terms of recall. Due to the longer answer length of LLaVA-Med and Quilt-LLaVA, these models are advantageous in the recall evaluation metric. The average length (word counts) of true answers is 18 and that number for LLaVA-Med and Quilt-LLaVA are 36 and 55 respectively, both of which are higher than BLIP-2 and our model (4 and 22 respectively). In addition, our precision exceeds LLaVA-Med’s by 10.25% and Quilt-LLaVA’s by 18.82 %, while our F1-score also surpasses LLaVA-Med’s by 6.53% and Quilt-LLaVA’s by 14.19%. In terms of a performance-cost ratio (the ratio of performance to the log of the number of parameters), CLOVER based on GPT-3.5 achieves the leading values for all metrics. CLOVER based on LLaMA-3.1 [35] and GPT-4o-mini also offer interesting advances. Overall, these findings indicate the strength of CLOVER to retain the high-level performance while minimizing costs.
@@
Table 1: Comparison with SOTA methods on PathVQA dataset. Bold and underline indicates the best and second best performances, respectively.